# Lookup multivariate Kolmogorov-Arnold Networks

### v1

@@ 3.3 Perspective on dimensions and spline orders
All the constructions introduced so far straightforwardly generalize to a higher-dimensional case. Evaluation of a single $d$- dimensional function parametrized by d-dimensional second-order B-splines would take $2^{d}$ add-multiplies, which stem from the summation of B-spline contributions residing on all the corners of the corresponding d-dimensional hypercube. Given that the number of such d-dimensional functions would be $d$ times smaller compared to the number of weights of a linear layer of the same shape, the inference FLOPs of the d-dimensional lmKAN layer would be $2^{d}/d$ of that of the linear layer with the same shape.
These slowdown factors are exactly identical, 2x, for one- and two-dimensional lmKANs, and start to grow for higher dimensions. Thus, we chose the two-dimensional version for the implementation of CUDA kernels, as this extension of the standard univariate KAN comes essentially for free.
If B-splines of order $k$ are employed for parametrization instead of the second-order ones described so far, the inference cost becomes $k^{d}/d$ of that of MLP. For more details, see the B-spline definition at Ref De Boor and De Boor (1978), and how they are used in KANs Liu et al. (2024). Increasing the B-splines order brings a few benefits, but they likely do not justify the associated increase in the computational cost.
The classical theorem about B-splines De Boor (1968) indicates that while the order of B-splines affects convergence rate, any spline order is sufficient to approximate any function arbitrarily close by increasing the resolution of the grid [^1] This is applicable, though, to functions on a bounded domain.. Given that the computational cost of spline look-up tables does not depend on the number of grid points, the grid resolution can be arbitrarily increased without any computational overhead at inference.
On top of that, increasing the B-spline order introduces additional smoothness of the model. Functions parametrized by B-splines of order $k$ belong to $C^{k-2}$, but in general not to $C^{k-1}$. The smoothness of second-order B-splines employed in this work matches that of ReLU Glorot et al. (2011), one of the most popular and successful activation functions, which is also continuous, but not continuously differentiable. Thus, it is questionable if the additional smoothness available through higher orders $k$ is necessary and would justify the associated computational overhead.
@@ 3.4 CUDA kernels
In this work, we implement CUDA kernels for efficient inference of 2D lmKANs on modern GPUs. Our kernels use the classic shared-memory tiling used in GEMM Volkov and Demmel (2008). In the following, all benchmarks run in full-precision float32.
With a $16\times 16$ tile, our implementation is $\sim 8\times$ slower than a dense linear layer with the same shape on an H100-SXM, irrespective of the grid resolution. The slowdown exceeds the $\sim 2\times$ FLOP-based estimate from Sec. 3.2, primarily because of the less coherent memory-access pattern of algorithm 1. Additionally, dense matrix multiplication has been the cornerstone of many computational pipelines and, thus, has enjoyed decades of thorough optimization.
@@
Reducing the tile to $8\times 8$ raises the slowdown to $\sim 9.5\times$ but lets us increase $G$ to 40, yielding $\approx\mathbf{88.5\times}$ better per-parameter efficiency.
@@ 3.5 Hessian regularization
When dealing with a similar problem, Xie et al. (2023) employed off-diagonal regularization based on a finite-difference scheme for the second derivative. One can put regularization terms as $\lambda\sum_{i}(p_{i}-p_{i+1})^{2}$ for first derivative, $\lambda\sum_{i}(p_{i}-2p_{i+1}+p_{i+2})^{2}$ for second, and so on. Such regularization schemes result in meaningful approximations after training, as illustrated in Fig. 4b.
We implemented CUDA kernels for 2D lmKANs, which express general high-dimensional mapping in terms of building blocks of two-dimensional functions. For 2D functions, we use off-diagonal regularization based on the squared Frobenius norm of the Hessian, a rotationally invariant measure of curvature in any direction, which is not to be confused with the Laplacian. Furthermore, our finite-differences schemes take into account that the grids, introduced in Sec. 3.1, are not uniform. More details are given in Appendix. C.
The Hessian of a function is zero if and only if the function is linear. Therefore, the use of a very strong Hessian-based regularization leads to linearization of the trained functions, enforcing them to converge to $f(x_{1},x_{2})=ax_{1}+bx_{2}+c$. This makes the whole lmKAN equivalent to an MLP of the same shape, modulus training dynamics. In other words, the Hessian regularization coefficient $\lambda$ can be used to smoothly adjust the lmKAN behavior between fully unconstrained lmKAN and MLP extremes. This observation that lmKANs with heavy Hessian regularization match non-regularized MLPs suggests that one should use a combination of the proposed regularization scheme with standard ones, such as L2 or dropout Srivastava et al. (2014), for the best results.
@@
Similarly to the original KANs Liu et al. (2024), we append a preconditioning term to splined functions, which in our case is linear. To further increase training stability, we employ a multi-staged fitting procedure, where the strength of the described Hessian regularization is initially set to a high value and then gradually decays. More details are available in Appendix E.
@@ 4 Experiments
It has been demonstrated so far that lmKANs can have significantly better inference cost per trainable parameter compared to linear layers in terms of both FLOPs and wall clock time on modern GPUs. The question is, however, whether this nominal efficiency translates to real-life performance. Do lmKANs indeed represent a better tradeoff between performance and inference cost?
In this section, we empirically compare the efficiency of lmKANs and MLPs across three settings: (i) approximating general high-dimensional functions, (ii) on a tabular-like dataset of randomly displaced methane configurations, and (iii) within CNN frameworks evaluated on CIFAR-10 and ImageNet. Across all tasks, lmKANs are consistently inference FLOPs — accuracy Pareto-optimal, with the largest gains on the methane dataset.
@@ 4.1 General function approximation
We fit both MLP and lmKAN students to approximate this ground-truth function and compare their performance. We use the same fully-connected backbone for both types of models with two hidden layers and varying hidden dimensions. Both MLPs and lmKANs use batch normalizations. We set affine=True for MLPs as it is the standard choice, and affine=False for lmKANs in accordance with static percentile grids introduced in Sec. 3.1. MLPs use ReLU activations, while lmKANs do not require any additional activation functions. We use $G=12$ for all the lmKAN models, as this was the optimal value found in the ablation study described below. Pseudo code for both the models is available in Fig. 5. Both students have two hidden layers, which is one more than both Cybenko Cybenko (1989) (the one for MLPs) and Kolmogorov-Arnold universal approximation theorems require. This setup, however, is more realistic, as MLPs with exactly one hidden layer are rarely used in practice.
@@
We fit all the models with the Adam optimizer Kingma and Ba (2014). For each step of stochastic gradient descent, we generate random inputs from the normal distribution, and compute the corresponding targets by evaluating the ground-truth teacher MLP. In such an infinite data regime, the final Mean-Squared Error (MSE) depends on how flexible the models are, and monotonically decreases with the hidden dimension for both MLPs and lmKANs.
The increase of the hidden dimension inevitably entails a higher computational cost at inference, and the question is which family of models represents a better tradeoff between accuracy and computational efficiency. Our findings are summarized in Fig. 6. The left column illustrates the final MSE of converged models depending on the hidden dimension, the middle column represents the Pareto front between the final MSE and FLOPS required at inference, and the last column contains the Pareto front between the final MSE and inference H100 wall-clock time.
In order to justify the claims that lmKAN are indeed more efficient, we converge baseline MLP-based models very tightly here, and in all similar experiments in this manuscript. For the MLP baseline, we have two lines - one with a full training budget and one with only half of it. The fact that these lines nearly coincide with each other demonstrates very tight convergence.
@@
Contrary to our initial expectations, the final MSE does not monotonously decrease with the grid resolution, having a distinct minimum at $G=12$. This is happening despite the fact that our setup effectively corresponds to an infinite data regime, effectively ruling out overfitting as an explanation. We hypothesize that the reason is that Kolmogorov-Arnold Networks are hard to converge for excessively fine grid resolution. Therefore, the optimal value of $G$ could depend on the training protocol. Appendix F.2 reinforces this supposition by providing the MSE vs. epoch number plot with several lines corresponding to different $G$ values. Additionally, note that FastKANs were found to display the same effect to an even greater degree, as detailed in Sec. 4.4.
@@ 4.2 Randomly displaced methane configurations
Our next step was to benchmark the performance of lmKANs on real data. Tabular datasets are the natural choice for feed forward fully-connected neural networks. Popular tabular datasets, such as the Titanic Kaggle or housing prices Harrison Jr and Rubinfeld (1978) ones, however, are not particularly convenient for this purpose. First, they are typically stochastic in nature - for instance, while it is possible to improve a guess on the survival based on the data available for the Titanic dataset, it is impossible to say for sure. Thus, even an arbitrarily large model fitted on arbitrarily many data points would have a non-zero limitation on the accuracy. In other words, the performance of a model translates into an error metric not so directly, making the comparisons between different models less illustrative. Second, these datasets are typically relatively small, making it challenging to sweep across a wide range of model scales to obtain a comprehensive picture of performance.
@@
Therefore, we chose the tabular-like dataset of randomly displaced methane configurations for the comparison. It consists of multiple off-equilibrium methane configurations, as illustrated in Fig. 8. The target is given by the corresponding quantum-mechanical energy Turney et al. (2012); Kohn and Sham (1965). Hydrogen atoms are placed around the carbon atom not as an ideal tetrahedron of an equilibrium methane molecule, but rather randomly, varying from instance to instance. Thus, the corresponding quantum-mechanical energies of such configurations are also different from each other. Machine learning models fitted on such datasets belong to the class of so-called machine learning interatomic potentials Behler and Parrinello (2007); Bartók et al. (2010). This dataset is sufficiently large for the comparisons, containing more than seven million configurations. Additionally, this dataset is deterministic - the geometry of the corresponding methane configuration completely determines the target (Formally, there can be a stochastic term due to the lack of complete convergence of ab initio computations for the quantum-mechanical energy, but it is negligible in practice).
@@
The target, the potential energy of the system, is invariant with respect to rotations and permutations of identical atoms [^2] Formally, there is an additional symmetry, inversion, but the corresponding group contains only two elements, thus it does not make much sense to treat it separately. We unite it with the group of rotations, and in the following, by rotation we mean proper or improper rotation. . Therefore, there are several viable representations of the methane molecules depending on how these symmetries are addressed:
@@
Cartesian Components: The simplest representation is just a collection of all the Cartesian components of all displacement vectors from the carbon atom to all the hydrogen atoms. Since each methane molecule contains $4$ hydrogen atoms, the total number of displacement vectors is $4$, and the total number of components is $12$. When using this representation, we simply concatenate all these components together and feed them to a fully connected MLP or lmKAN whose input dimension is $12$. This representation is not invariant with respect to both rotations and permutations; thus, we use the corresponding augmentations during training. We randomly permute hydrogen atoms and rotate each molecule whenever we sample a minibatch from the training subset for each step of stochastic gradient descent.
@@
Cartesian Components Polynomials: We compute power sum symmetric polynomials on top of the Cartesian components of the displacements vectors: $P_{\alpha_{x},\alpha_{y},\alpha_{z}}=\sum_{i=1}^{i=4}x_{i}^{\alpha_{x}}y_{i}^{\alpha_{y}}z_{i}^{\alpha_{z}}$ for non-negative integer $\alpha_{x}+\alpha_{y}+\alpha_{z}\leq 4$. The total number of such symmetric polynomials is $34$ (excluding trivial $P_{0,0,0}$). This representation is invariant with respect to permutations but not with respect to rotations. Thus, during training we use only rotational augmentations.
Distances Polynomials: The final representation is a collection of non-trivial symmetric polynomials on top of the interatomic distances, constructed similarly as in Allen et al. (2021). The total number of such polynomials is $31$, and their exact formulas are given in Appendix. F.3. This representation is invariant with respect to both rotations and permutations. Thus, we do not use any augmentations during training for this representation.
@@
For each representation, we fit the same families of MLP and lmKAN models as in the previous section. The result is given in Fig. 9. For this dataset we use $G=28$, the optimal value we found in ablation studies. Similarly to the previous experiment, we demonstrate tight convergence of the baseline MLP models by providing three lines corresponding to full, half, and quarter of the training budget, respectively. Overall, when compared to domain-specific architectures, typically given by GNNs Zhang et al. (2021) and/or transformers Pozdnyakov and Ceriotti (2023), the introduced feed-forward fully-connected models occupy a non-overlapping part of the Pareto frontier — they are less accurate, but also orders of magnitude faster.
The figure illustrates that lmKANs consistently outperform MLPs across all modalities. Furthermore, the performance improvement is much larger compared to our previous experiment. At the same accuracy level, lmKANs require up to many dozen times (or even more for the Distances modality) less inference FLOPS, which results in more than a 10× improvement of the inference H100 wall clock time.
Furthermore, Fig. 9 provides early indications that lmKANs sometimes can be more accurate in the limit of large scale, that is, to have better generalizability. The second row of the figure, corresponding to the Distances modality, illustrates that the rate of improvement of MLP models becomes very slow, and it is questionable if this family of models would ever surpass the accuracy achieved by lmKAN models at any scale.
On the other hand, depending on the nature of the data, raw lmKANs, without the Hessian regularization we proposed in Sec. 3.5, can be more prone to overfitting. This happens for the Distances Polynomials modality as the last row of Fig. 9 illustrates. This modality incorporates all the symmetries into the representation and does not involve any sort of augmentations. Therefore, it is likely that the generalization problem we outlined in Sec. 3.5 takes place for this fitting setup. As the green line of the fourth row of Fig. 9 illustrates, the Hessian regularization is sufficient to overcome the overfitting. Properly regularized lmKANs were found to outperform the MLPs and be Pareto optimal from the point of view of both inference FLOPS and inference H100 wall clock time.
@@
In the introduction, we briefly outlined that high-dimensional linear mappings are the primary building blocks in most architectures, not only in feed-forward fully connected neural networks. Convolutional Neural Networks are no exception.
A standard computer-vision two-dimensional convolution with kernel size $k\times k$ is parametrized by linear mapping $\mathbb{R}^{k^{2}C_{in}}\rightarrow\mathbb{R}^{C_{out}}$, where $C_{in}$ and $C_{out}$ are numbers of input and output channels respectively. Since Kolmogorov-Arnold layers can be used as a general substitute for high-dimensional linear mappings, one can construct a KAN-based convolution neural network well suited for image processing, as was done, e.g, in Bodner et al. (2024). In this section, we compare the performance of lmKAN- and MLP-based CNNs on the CIFAR-10 dataset Krizhevsky et al. (2009) and ImageNet Deng et al. (2009) datasets.

### v2

@@ 3.3 Perspective on dimensions and spline orders
All the constructions introduced so far straightforwardly generalize to a higher-dimensional case. Evaluation of a single d-dimensional function parametrized by d-dimensional second-order B-splines would take $2^{d}$ multiply-adds, which stem from the summation of B-spline contributions residing on all the corners of the corresponding d-dimensional hypercube. Given that the number of such d-dimensional functions would be $d$ times smaller compared to the number of weights of a linear layer of the same shape, the inference FLOPs of the d-dimensional lmKAN layer would be $2^{d}/d$ of that of the linear layer with the same shape.
These slowdown factors are exactly identical, 2×, for one- and two-dimensional lmKANs, and start to grow for higher dimensions. Thus, we chose the two-dimensional version for the implementation of CUDA kernels, as this extension of the standard univariate KAN comes essentially for free.
If B-splines of order $k$ are employed for parametrization instead of the second-order ones described so far, the inference cost becomes $k^{d}/d$ of that of a linear layer. For more details, see the B-spline definition at De Boor and De Boor [32], and how they are used in KANs [11]. Increasing the B-spline order brings a few benefits, but they likely do not justify the associated increase in the computational cost.
The classical theorem about B-splines [33] indicates that while the order of B-splines affects the convergence rate, any spline order is sufficient to approximate any function arbitrarily closely by increasing the resolution of the grid [^1] This is applicable, though, to functions on a bounded domain.. Given that the computational cost of spline look-up tables does not depend on the number of grid points, the grid resolution can be arbitrarily increased without any computational overhead at inference.
On top of that, increasing the B-spline order introduces additional smoothness of the model. Functions parametrized by B-splines of order $k$ belong to $C^{k-2}$, but in general not to $C^{k-1}$. The smoothness of second-order B-splines employed in this work matches that of ReLU [34], one of the most popular and successful activation functions, which is also continuous, but not continuously differentiable. Thus, it is questionable if the additional smoothness available through higher orders $k$ is necessary and would justify the associated computational overhead.
@@ 3.4 CUDA kernels
In this work, we implement CUDA kernels for efficient inference of 2D lmKANs on modern GPUs. Our kernels use the classic shared-memory tiling used in GEMM [35]. In the following, all benchmarks run in full-precision float32.
With a $16\times 16$ tile, our implementation is $\sim 8\times$ slower than a dense linear layer with the same shape on an H100-SXM, irrespective of the grid resolution. The slowdown exceeds the $\sim 2\times$ FLOPs-based estimate from Sec. 3.2, primarily because of the less coherent memory-access pattern of algorithm 1. Additionally, dense matrix multiplication has been the cornerstone of many computational pipelines and, thus, has enjoyed decades of thorough optimization.
@@
Reducing the tile to $8\times 8$ raises the slowdown to $\sim 9.5\times$ but lets us increase $G$ to 40, yielding $\sim\mathbf{88.5\times}$ better per-parameter efficiency.
@@ 3.5 Hessian regularization
When dealing with a similar problem, Xie et al. [36] employed off-diagonal regularization based on a finite-difference scheme for the second derivative. One can put regularization terms as $\lambda\sum_{i}(p_{i}-p_{i+1})^{2}$ for first derivative, $\lambda\sum_{i}(p_{i}-2p_{i+1}+p_{i+2})^{2}$ for second, and so on. Such regularization schemes result in meaningful approximations after training, as illustrated in Fig. 4b.
We implemented CUDA kernels for 2D lmKANs, which express general high-dimensional mapping in terms of building blocks of two-dimensional functions. For 2D functions, we use off-diagonal regularization based on the squared Frobenius norm of the Hessian, a rotationally invariant measure of curvature in any direction, which is not to be confused with the Laplacian. Furthermore, our finite-differences schemes take into account that the grids, introduced in Sec. 3.1, are not uniform. More details are given in Appendix C.
The Hessian of a function is zero if and only if the function is linear. Therefore, the use of a very strong Hessian-based regularization leads to linearization of the trained functions, enforcing them to converge to $f(x_{1},x_{2})=ax_{1}+bx_{2}+c$. This makes the whole lmKAN equivalent to an MLP of the same shape, modulo training dynamics. In other words, the Hessian regularization coefficient $\lambda$ can be used to smoothly adjust the lmKAN behavior between fully unconstrained lmKAN and MLP extremes. This observation—that lmKANs with heavy Hessian regularization match non-regularized MLPs—suggests that one should use a combination of the proposed regularization scheme with standard ones, such as L2 or dropout [37], for the best results.
@@
Similarly to the original KANs [11], we append a preconditioning term to splined functions, which in our case is linear. To further increase training stability, we employ a multi-staged fitting procedure, where the strength of the described Hessian regularization is initially set to a high value and then gradually decays. More details are available in Appendix E.
@@ 4 Experiments
We have demonstrated so far that lmKANs can have significantly better inference cost per trainable parameter compared to linear layers in terms of both FLOPs and wall-clock time on modern GPUs. The question is, however, whether this nominal efficiency translates to real-life performance. Do lmKANs indeed represent a better trade-off between performance and inference cost?
In this section, we empirically compare the efficiency of lmKANs and MLPs across the following settings: (i) approximating general high-dimensional functions, (ii) on a tabular-like dataset of randomly displaced methane configurations, and (iii) within CNN frameworks evaluated on CIFAR-10 and ImageNet. Across all experiments, we use identical macro-architectural backbones for lmKANs and MLPs. Overall, to obtain a comprehensive picture of the performance, we prioritized the diversity of the setups over a very large scale or the architectural complexity of a particular backbone. We found that lmKANs are consistently inference FLOPs — accuracy Pareto-optimal, with the largest gains on the methane dataset. Finally, we compare lmKANs with FastKANs.
@@ 4.1 General function approximation
We fit both MLP and lmKAN students to approximate this ground-truth function and compare their performance. We use the same fully connected backbone for both types of models with two hidden layers and varying hidden dimensions. Both MLPs and lmKANs use batch normalizations. We set affine=True for MLPs as it is the standard choice, and affine=False for lmKANs in accordance with static percentile grids introduced in Sec. 3.1. MLPs use ReLU activations, while lmKANs do not require any additional activation functions. We use $G=12$ for all the lmKAN models, as this was the optimal value found in the ablation study described below. Pseudocode for both models is available in Fig. 5. Both students have two hidden layers, which is one more than both Cybenko [38] (the one for MLPs) and Kolmogorov-Arnold universal approximation theorems require. This setup, however, is more realistic, as MLPs with exactly one hidden layer are rarely used in practice.
@@
We fit all the models with the Adam optimizer [39]. For each step of stochastic gradient descent, we generate random inputs from the normal distribution, and compute the corresponding targets by evaluating the ground-truth teacher MLP. In such an infinite data regime, the final Mean-Squared Error (MSE) depends on how flexible the models are, and monotonically decreases with the hidden dimension for both MLPs and lmKANs.
The increase of the hidden dimension inevitably entails a higher computational cost at inference, and the question is which family of models represents a better trade-off between accuracy and computational efficiency. Our findings are summarized in Fig. 6. The left column illustrates the final MSE of converged models depending on the hidden dimension, the middle column represents the Pareto front between the final MSE and FLOPs required at inference, and the last column contains the Pareto front between the final MSE and inference H100 wall-clock time.
In order to justify the claims that lmKANs are indeed more efficient, we converge baseline MLP-based models very tightly here, and in all similar experiments in this manuscript. For the MLP baseline, we have two lines - one with a full training budget and one with only half of it. The fact that these lines nearly coincide with each other demonstrates very tight convergence.
@@
Contrary to our initial expectations, the final MSE does not monotonically decrease with the grid resolution, having a distinct minimum at $G=12$. This is happening despite the fact that our setup effectively corresponds to an infinite data regime, effectively ruling out overfitting as an explanation. We hypothesize that the reason is that Kolmogorov-Arnold Networks are hard to converge for excessively fine grid resolution. Therefore, the optimal value of $G$ could depend on the training protocol. Appendix F.2 reinforces this supposition by providing the MSE vs. epoch number plot with several lines corresponding to different $G$ values. Additionally, note that FastKANs were found to display the same effect to an even greater degree, as detailed in Sec. 4.4.
@@ 4.2 Randomly displaced methane configurations
Our next step was to benchmark the performance of lmKANs on real data. Tabular datasets are the natural choice for feedforward fully connected neural networks. Popular tabular datasets, such as the Titanic [40] or housing prices [41], however, are not particularly convenient for this purpose. First, they are typically stochastic in nature - for instance, while it is possible to improve a guess on the survival based on the data available for the Titanic dataset, it is impossible to say for sure. Thus, even an arbitrarily large model fitted on arbitrarily many data points would have a non-zero limitation on the accuracy. In other words, the performance of a model translates into an error metric not so directly, making the comparisons between different models less illustrative. Second, these datasets are typically relatively small, making it challenging to sweep across a wide range of model scales to obtain a comprehensive picture of performance.
@@
Therefore, we chose the tabular-like dataset of randomly displaced methane configurations [42] for the comparison. It consists of multiple off-equilibrium methane configurations, as illustrated in Fig. 8. The target is given by the corresponding quantum-mechanical energy [43, 44]. Hydrogen atoms are placed around the carbon atom not as an ideal tetrahedron of an equilibrium methane molecule, but rather randomly, varying from instance to instance. Thus, the corresponding quantum-mechanical energies of such configurations are also different from each other. Machine learning models fitted on such datasets belong to the class of so-called machine learning interatomic potentials [45, 46]. This dataset is sufficiently large for the comparisons, containing more than seven million configurations. Additionally, this dataset is deterministic - the geometry of the corresponding methane configuration completely determines the target (Formally, there can be a stochastic term due to the lack of complete convergence of ab initio computations for the quantum-mechanical energy, but it is negligible in practice).
@@
The target, the potential energy of the system, is invariant with respect to rotations and permutations of identical atoms [^2] Formally, there is an additional symmetry, inversion, but the corresponding group contains only two elements, thus it does not make much sense to treat it separately. We treat it as part of the rotation group, and in the following, by rotation we mean proper or improper rotation. . Therefore, there are several viable representations of the methane molecules depending on how these symmetries are addressed:
@@
Cartesian Components: The simplest representation is a collection of all the Cartesian components of all displacement vectors from the carbon atom to all the hydrogen atoms. Since each methane molecule contains $4$ hydrogen atoms, the total number of displacement vectors is $4$, and the total number of components is $12$. When using this representation, we simply concatenate all these components together and feed them to a fully connected MLP or lmKAN whose input dimension is $12$. This representation is not invariant with respect to both rotations and permutations; thus, we use the corresponding augmentations during training. We randomly permute hydrogen atoms and rotate each molecule whenever we sample a minibatch from the training subset for each step of stochastic gradient descent.
@@
Cartesian Components Polynomials: We compute power sum symmetric polynomials on top of the Cartesian components of the displacement vectors: $P_{\alpha_{x},\alpha_{y},\alpha_{z}}=\sum_{i=1}^{i=4}x_{i}^{\alpha_{x}}y_{i}^{\alpha_{y}}z_{i}^{\alpha_{z}}$ for non-negative integer $\alpha_{x}+\alpha_{y}+\alpha_{z}\leq 4$. The total number of such symmetric polynomials is $34$ (excluding trivial $P_{0,0,0}$). This representation is invariant with respect to permutations but not with respect to rotations. Thus, during training we use only rotational augmentations.
Distances Polynomials: The final representation is a collection of non-trivial symmetric polynomials on top of the interatomic distances, constructed similarly as in Allen et al. [47]. The total number of such polynomials is $31$, and their exact formulas are given in Appendix F.3. This representation is invariant with respect to both rotations and permutations. Thus, we do not use any augmentations during training for this representation.
@@
For each representation, we fit the same families of MLP and lmKAN models as in the previous section. The result is given in Fig. 9. For this dataset, we use $G=28$, the optimal value we found in ablation studies. Similarly to the previous experiment, we demonstrate tight convergence of the baseline MLP models by providing three lines corresponding to full, half, and quarter of the training budget, respectively. Overall, when compared to domain-specific architectures, typically given by GNNs [48] and/or transformers [49], the introduced feedforward fully connected models occupy a non-overlapping part of the Pareto frontier — they are less accurate, but also orders of magnitude faster.
The figure illustrates that lmKANs consistently outperform MLPs across all modalities. Furthermore, the performance improvement is much larger compared to our previous experiment. At the same accuracy level, lmKANs require dozens of times (or even up to one thousand for the Distances modality) fewer inference FLOPs, which results in more than a 10× improvement of the inference H100 wall-clock time.
On top of that, Fig. 9 provides early indications that lmKANs sometimes can be more accurate in the limit of large scale, that is, to have better generalizability. The second row of the figure, corresponding to the Distances modality, illustrates that the rate of improvement of MLP models becomes very slow, and it is questionable if this family of models would ever surpass the accuracy achieved by lmKAN models at any scale.
On the other hand, depending on the nature of the data, raw lmKANs, without the Hessian regularization we proposed in Sec. 3.5, can be more prone to overfitting. This happens for the Distances Polynomials modality as the last row of Fig. 9 illustrates. This modality incorporates all the symmetries into the representation and does not involve any sort of augmentations. Therefore, it is likely that the generalization problem we outlined in Sec. 3.5 takes place for this fitting setup. As the green line of the fourth row of Fig. 9 illustrates, the Hessian regularization is sufficient to overcome the overfitting. Properly regularized lmKANs were found to outperform the MLPs and be Pareto optimal from the point of view of both inference FLOPs and inference H100 wall-clock time.
@@
In the introduction, we briefly outlined that high-dimensional linear mappings are the primary building blocks in most architectures, not only in feedforward fully connected neural networks. Convolutional Neural Networks are no exception.
A standard computer-vision two-dimensional convolution with kernel size $k\times k$ is parametrized by linear mapping $\mathbb{R}^{k^{2}C_{in}}\rightarrow\mathbb{R}^{C_{out}}$, where $C_{in}$ and $C_{out}$ are numbers of input and output channels, respectively. Since Kolmogorov-Arnold layers can be used as a general substitute for high-dimensional linear mappings, one can construct a KAN-based convolutional neural network well suited for image processing, as was done, e.g, in Bodner et al. [50]. In this section, we compare the performance of lmKAN- and MLP-based CNNs on the CIFAR-10 [51] and ImageNet [52] datasets.